OpenStudy (anonymous):

Solve for t. (1/2)t - 6 = -11

OpenStudy (anonymous):

t=-10

OpenStudy (anonymous):

Any work

OpenStudy (anonymous):

(1/2)t=-11+6=-5 t=-5/1/2=-5*2=-10